OpenStudy (anonymous):

in the expansion of (x+y)^15 a certain term contains x^10. What is the exponent of y in this term? a.8 b.9 c.11 d.5

OpenStudy (asnaseer):

in the binomial expansion of:\[(x+y)^{15}\]you know the first term will be \(x^{15}\) and the last term will be \(y^{15}\). every term in between will be of the form:\[Cx^ay^b\]where C is some constant and a+b=15

OpenStudy (asnaseer):

can you work out the answer now?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

think about it and re-read what I have written. remember you are told the term in question has \(x^{10}\) in it. so using my notation you know that a=10 therefore: a + b = 15 10 + b = 15 therefore b=?

OpenStudy (anonymous):

5

OpenStudy (asnaseer):

correct!

OpenStudy (anonymous):

thank you :)

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

ok and i have one more question

OpenStudy (asnaseer):

just type each question in the list to the left